OpenStudy (natalie1234):

HELP WILL GIVE MEDAL! Find a number that is between 2/7 and 4/11

OpenStudy (anonymous):

or.. just divide the numbers and pick a number in between... if you want a decimal number

OpenStudy (anonymous):

Divide 2 by 7 and 4 by 11.. You get.. .28 and .36 Now pick a number in between

OpenStudy (anonymous):

2/7=22/77 4/11=28/77 you can chose 23/77 or 24/77 or 25/77 or 26/77 or 27/77

OpenStudy (anonymous):

Or what @d3v13 said

OpenStudy (natalie1234):

okay thanks guys for the help (: i got my answer